# Cognitive Bias in Decision-Making with LLMs

### v2

@@ Abstract
Large language models (LLMs) offer significant potential as tools to support an expanding range of decision-making tasks. Given their training on human (created) data, LLMs have been shown to inherit societal biases against protected groups, as well as be subject to bias functionally resembling cognitive bias. Human-like bias can impede fair and explainable decisions made with LLM assistance. Our work introduces BiasBuster, a framework designed to uncover, evaluate, and mitigate cognitive bias in LLMs, particularly in high-stakes decision-making tasks. Inspired by prior research in psychology and cognitive science, we develop a dataset containing 16,800 prompts to evaluate different cognitive biases (e.g., prompt-induced, sequential, inherent). We test various bias mitigation strategies, amidst proposing a novel method utilising LLMs to debias their own prompts. Our analysis provides a comprehensive picture of the presence and effects of cognitive bias across commercial and open-source models. We demonstrate that our self-help debiasing effectively mitigates model answers that display patterns akin to human cognitive bias without having to manually craft examples for each bias.
@@
Jessica Echterhoff                        Yao Liu                        Abeer Alessa University of California, San Diego {jechterh, yal004, aalessa, jmcauley, zehe}@ucsd.edu                        Julian McAuley                        Zexue He
@@ 1 Introduction
LLMs exhibit strong performance across multiple tasks Albrecht et al. (2022), such as summarizing documents Wang et al. (2023), answering math questions Imani et al. (2023) or chat-support Lee et al. (2023). These capabilities lead humans to increasingly use LLMs for support or advice in their day-to-day decisions Rastogi et al. (2023); Li et al. (2022). However, models suffer from various algorithmic biases, requiring procedures to evaluate and mitigate bias Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021). In addition to societal bias, LLMs can show answer patterns similar to human-like cognitive bias, which can implicitly mislead a user’s decision-making Schramowski et al. (2022). Cognitive bias refers to a systematic pattern of deviation from norms of rationality in judgment, where individuals create their own “subjective reality” from their perception of the input  Haselton et al. (2015); Kahneman et al. (1982), and leads to inconsistent decision-making. Cognitive bias arises in human decision-making as well as human-ML interaction Bertrand et al. (2022). Although language models do not possess cognition, they might show signs of bias that functionally resemble human cognitive bias. Hence, when LLMs aid humans in high-stakes decision-making, such as evaluating individuals, these models must be properly audited Rastogi et al. (2023).
@@
Figure 1: In high-stakes decision-making, BiasBuster assesses model outputs for patterns similar to human cognitive biases and tests various bias mitigation techniques.
@@
Different from societal bias where behavior is influenced by social and cultural background, cognitive bias arises from the information processing mechanisms in human decision-making procedures, often influenced by the setup of the task. Cognitive bias is often not directly visible and hence difficult to detect. Our work introduces a novel approach to quantify and mitigate patterns akin to human cognitive bias in LLMs using cognitive bias-aware prompting techniques.
@@
Our work proposes BiasBuster (Figure 1), a systematic framework that encapsulates quantitative evaluation and automatic mitigation procedures for cognitive bias. To evaluate human-like cognitive bias in LLMs, BiasBuster provides an extended set of testing prompts for a variety of biases which are developed in accordance with cognitive science experiments, but aligned for LLMs. We develop metrics to evaluate the response of large language models to prompts classified as “biased” and “neutral” with respect to cognitive bias for humans. BiasBuster compares different debiasing strategies, some shown to be effective on humans, in zero-shot and few-shot prompting. To minimize manual effort in prompt creation, we propose a novel prompting strategy where a language model debiases its prompts and helps itself to be less subject to bias (we call it self-help). BiasBuster provides a thorough evaluation of different debiasing methods, enabling practitioners to effectively address bias.
@@
To avoid cross-contamination with existing data that a model might have been trained on, BiasBuster provides novel prompts for a high-stakes decision-making scenario – student admissions for a college program, where we generate and provide sets of cognitive bias testing and debiasing prompts. These testing prompts quantitatively evaluate various patterns for cognitive biases in terms of LLM self-consistency and decision confidence. The debiasing prompts assess the utility of various mitigation techniques, specifically focusing on the ability of LLMs to de-bias their prompts.
@@ 2.1 Bias in Large Language Models
Many different social biases Liang et al. (2021) and biases related to reasoning and decision-making Itzhak et al. (2023) have been detected in LLMs. Viswanath and Zhang (2023) propose a toolkit for evaluating social biases in LLMs, including evaluation metrics for detecting social biases Ribeiro et al. (2020) perform a test comprising a small set of neutral sentences with simple adjectives, label preserving perturbations to check if the behavior of the LLM differs, and test adding a sentiment to the template to check if the model predicts the opposite sentiment Ribeiro et al. (2020). Compared to their work, which focuses on the extent of biased decisions that are made towards protected groups, our work provides insight into decision patterns akin to human cognitive bias where we analyze systematic flaws of language models during a decision-making procedure.
Existing evaluation metrics for societal bias are often based on word embeddings Bolukbasi et al. (2016); Papakyriakopoulos et al. (2020); Viswanath and Zhang (2023), which is not directly applicable for evaluation of decision patterns akin to human cognitive bias. Cognitive bias is not necessarily embedded in specific tokens, but can be reflected in the entire current Tversky and Kahneman (1981) or previous context Echterhoff et al. (2022).
@@ 2.2 Patterns of Cognitive Bias in LLMs
To address the lack of evaluation metrics for cognitive bias, Lin and Ng (2023) propose metrics for availability and framing bias using dummy inputs and paraphrasing for classification tasks. In comparison, our work focuses on generative tasks. Previous work has tackled individual cognitive biases such as representativeness, base rate neglect, etc. Talboy and Fuller (2023). However, they evaluate the biases as one-off questions and answers, which require further investigation on generalization. Jones and Steinhardt (2022) test for systematic qualitative errors of LLM responses with human cognitive bias using coding prompts. Cognitive bias can influence many scenarios, but is especially important for high-stakes decisions in human-AI collaboration, as humans might be influenced by the decision assistance that LLMs provide.
@@ 3 Testing for Patterns of Cognitive Bias in LLMs
We classify cognitive bias into three categories. Prompt-based cognitive bias describes a subset of cognitive bias mainly introduced through a user-prompt. Inherent bias is instilled in the model through the training data. Inherent bias is not necessarily obvious in the prompt but by the general task itself. Sequential bias is induced by previous model answers.
@@
Anchoring Imagine you are an admissions officer. Typically 30% of students are admitted to the program. Would you admit or reject the following student to the program? The student is from {country} and studied {major} at {school} with a GPA of {gpa}, […]
Status Quo/ Primacy Status Quo:“You are aiming to admit a student to your lab. You have worked with student A during a summer internship. Here are student features to help your decision: a) Student A: Currently interning in your lab. They have a background in {background}. They are experienced in {experience}, but need to develop {skill} skills. … d) Student D: Comes from {university} with strong skills in {skill}. They specialize in {topic}, but lack experience in {experience}. Select option a, b, c or d.” Neutral:“You are aiming to admit a student to your university lab. Here are student features to help your decision: a) Student A: They have a background in {background}. They are experienced in {experience}, but need to develop {skill} skills. … Select option a, b, c or d.”
@@
We design experiments to test decision patterns similar to human cognitive bias, based on definitions from human cognitive bias literature. Our work aims to align all bias groups (prompt-based, sequential, inherent) as much as possible with the same evaluation metrics. However, the detection of individual biases in each group has to be tackled with separate metrics to be able to account for the nuances of the bias group. For all biases, we strive to find a metric of “consistency”. Compared to evaluating decision pattterns on human participants, LLMs have the distinct advantage of being testable under various study conditions through repeated prompting.
@@ 3.1 Sequential Bias
Anchoring bias describes the human tendency to change perception based on an anchor Kahneman et al. (1982). We follow the setup of (Echterhoff et al., 2022), in which decision-makers are influenced (anchored) by their own recent decisions. This setup aims to evaluate bias in sequential setups, compared to one-off prompt-based setups, which we discuss in the next section.
@@
To analyse the influence of previous decisions in language models, we ask the model to take the role of an admissions officer deciding which student to admit to a college study program. We create synthetic student profiles and show them to the language model in one conversation by always adding the previous student and the model’s previous decision to the context. We perturb different student sets such that the same set of students is exposed to the model in different orders, to observe if LLMs make different decisions for the same students. We show examples of our templates in Table 1.
@@ 3.2 Prompt-Based Cognitive Bias
Status quo bias is a cognitive bias that refers to the tendency of people to prefer and choose the current state of affairs or the existing situation over change or alternative options Samuelson and Zeckhauser (1988). Given a set of questions that differ in their content by providing a default option in the status quo, a biased question can be compared to the same prompt without status quo information (neutral condition). Questions always provide different options to choose from. We take inspiration from Samuelson and Zeckhauser (1988) which bias the user with a status quo option with respect to car brands and investment options to choose from. Given e.g. a current car brand they drive or a current investment, users then have to make a decision to switch their car or investment or keep the status quo.
@@
We develop a template for testing if a model shows decision patterns similar to status quo bias between a neutral question, which has no information on current status, and a status quo question for the student admissions setup. In this case, we ask for a student to be admitted to a research lab given student features, and provide four options to choose from. We define the status quo to be “having worked with student X in a summer internship before”. Other parts of the question and the student options remain the same. From a pool of 16 student profiles, we choose 4 to be displayed at a time and show each student at each position to evaluate if some options are chosen disproportionally.
@@
In the status quo experiment, we have a single-choice problem setup, where for each question we can select exactly one option. As all students appear at each position for each student set, the distribution of chosen answers should be uniform. We measure if any option (A,B,C,D) is chosen more often than others. A model would suffer from status quo bias if the default option is chosen more often than other options, so if $\frac{n_{SQ}}{n}>>0.25$ for the number of times the status quo option was chosen ($n_{SQ}$) over all decisions $n$.
@@
To analyse the difference in admissions or rejection behavior, we observe the admissions rate $\frac{1}{n}\sum_{i=0}^{n}d_{i}$ for admission decisions where $d_{i}\in\{0,1\}$ for rejection/admission of a student for all students $i=[0,...,n]$, which should not be affected by the framing of the question.
@@ 3.3 Inherent Cognitive Bias
Bias Number of Prompts
Anchoring 5425
Status Quo/Primacy 1008
Framing 2000
Group Attribution 1000
Table 2: Number of prompt instances in our dataset per cognitive bias.
@@ 3.4 BiasBuster Prompt Dataset
In total, we provide a dataset that can be used to test the LLM on patterns akin to human cognitive bias in over $16,800$ individual decisions. We show the size of each bias dataset in Table 2. For all our prompts, we use the English language. We publish our dataset on Huggingface.
@@ 4 Mitigating Cognitive Bias in LLMs
There are different approaches to mitigating decision patterns similar to human cognitive bias in LLMs. We group these approaches into zero-shot approaches, which can give additional information about the existence of cognitive bias without giving any examples, few shot approaches which can give examples of specific desired or undesired behavior and self-mitigation approaches, which use the model to debias itself (Figure 2).
@@ 4.2 Few-Shot-Mitigation
Figure 2: Overview of different mitigation techniques and comparison to our self-help setup, which is tasked to de-bias its prompts. We give an example for status quo bias, where the bias-inducing part of the prompt (in red) is removed by self-help.
@@
For group attribution, we show the same student twice, once as female as male, and ask the model answers to be the same. For framing, we show an example of the same student in different framing and ask the model to give the same admission outcome. For status quo, we show an example where the current student is not the most suitable candidate, but is still selected. For anchoring, we show two different orders of the same students with different answers for the individuals (Table 6).
@@
We show examples for counterfactual and contrastive mitigation for each bias in the Appendix in Table 6.

### v3

@@ Abstract
Large language models (LLMs) offer significant potential as tools to support an expanding range of decision-making tasks. Given their training on human (created) data, LLMs have been shown to inherit societal biases against protected groups, as well as be subject to bias functionally resembling cognitive bias. Human-like bias can impede fair and explainable decisions made with LLM assistance. Our work introduces BiasBuster, a framework designed to uncover, evaluate, and mitigate cognitive bias in LLMs, particularly in high-stakes decision-making tasks. Inspired by prior research in psychology and cognitive science, we develop a dataset containing 13,465 prompts to evaluate LLM decisions on different cognitive biases (e.g., prompt-induced, sequential, inherent) [^1] https://huggingface.co/datasets/jecht/cognitive_bias. We test various bias mitigation strategies, while proposing a novel method utilizing LLMs to debias their own human-like cognitive bias within prompts. Our analysis provides a comprehensive picture of the presence and effects of cognitive bias across commercial and open-source models. We demonstrate that our selfhelp debiasing effectively mitigates model answers that display patterns akin to human cognitive bias without having to manually craft examples for each bias.
@@
Jessica Echterhoffthanks: Corresponding authors: jechterh@ucsd.edu,    zehe@ucsd.edu 1, Yao Liu1, Abeer Alessa1, Julian McAuley1, Zexue He 1,2 1University of California, San Diego 2MIT-IBM Watson AI Lab {jechterh, yal004, aalessa, jmcauley, zehe}@ucsd.edu
@@ 1 Introduction
LLMs exhibit strong performance across multiple tasks Albrecht et al. (2022), such as summarizing documents Wang et al. (2023), answering math questions Imani et al. (2023) or chat-support Lee et al. (2023). These capabilities lead humans to increasingly use LLMs for support or advice in their day-to-day decisions Rastogi et al. (2023); Li et al. (2022). However, models suffer from various algorithmic biases, requiring procedures to evaluate and mitigate bias Zhao et al. (2018); Nadeem et al. (2020); Liang et al. (2021); He et al. (2021). In addition to societal bias, LLMs can show answer patterns similar to human-like cognitive bias, which can implicitly mislead a user’s decision-making Schramowski et al. (2022). Cognitive bias refers to a systematic pattern of deviation from norms of rationality in judgment, where individuals create their own “subjective reality” from their perception of the input  Haselton et al. (2015); Kahneman et al. (1982), and leads to inconsistent decision-making. Cognitive bias arises in human decision-making as well as human-ML interaction Bertrand et al. (2022). Although language models do not possess cognition, they might show signs of bias that functionally resemble human cognitive bias. Hence, when LLMs aid humans in decision-making, such as evaluating individuals, these models must be properly audited Rastogi et al. (2023).
@@
Figure 1: BiasBuster assesses model outputs for patterns similar to human cognitive biases and tests various bias mitigation techniques.
@@
Cognitive and social biases are highly connected. Cognitive biases are systematic tendencies leading to error – such as the tendency to interpret information in a way that confirms and reinforces pre-existing beliefs and opinions. Connected to these are social biases, formed automatically by impressions of people, based on the social group that they are a member of (Commission et al., 2021). Different from societal bias where behavior is influenced by social and cultural background, cognitive bias arises from the information processing mechanisms in human decision-making procedures, often influenced by the setup of the task Tversky and Kahneman (1974). Cognitive bias is often not directly visible and hence difficult to detect. Our work introduces a novel approach to quantifying and mitigating patterns akin to human cognitive bias in LLMs using cognitive bias-aware prompting techniques.
@@
Our work proposes BiasBuster (Figure 1), a systematic framework that encapsulates quantitative evaluation and automatic mitigation procedures for human-like cognitive bias. To evaluate human-like cognitive bias in LLMs, BiasBuster provides an extended set of testing prompts for a variety of biases which are developed in accordance with cognitive science experiments, but aligned for LLMs. We create metrics to assess how large language models respond to prompts categorized as either “biased” or “neutral” in relation to human-like cognitive biases. BiasBuster compares different debiasing strategies (some shown to also be effective on humans) in zero-shot and few-shot prompting. To minimize manual effort in prompt creation, we propose a novel prompting strategy where a language model debiases its prompts and helps itself to be less subject to bias (we call it selfhelp). BiasBuster provides a thorough evaluation of different debiasing methods, enabling practitioners to effectively address bias.
@@
To avoid cross-contamination with existing data that a model might have been trained on, BiasBuster provides novel prompts for a high-stakes decision-making scenario – student admissions for a college program, where we generate and provide sets of cognitive bias testing prompts and debiased prompts. These testing prompts quantitatively evaluate various patterns for cognitive biases, focusing on LLM self-consistency and decision confidence. The debiased prompts assess the utility of various mitigation techniques, specifically focusing on the ability of LLMs to debias their prompts.
@@ 2.1 Bias in Large Language Models
Many different social biases Liang et al. (2021) and biases related to reasoning and decision-making Itzhak et al. (2023) have been detected in LLMs (e.g. gender bias Kotek et al. (2023); Vig et al. (2020); Zhao et al. (2018), religious bias Abid et al. (2021), stereotype bias Nadeem et al. (2020), occupational bias Kirk et al. (2021), sentiment bias Huang et al. (2019) or bias against disabled individuals Venkit et al. (2022)). Previous work typically treats one bias at a time, which makes a generalized evaluation difficult. Viswanath and Zhang (2023) propose a toolkit for evaluating social biases in LLMs, including evaluation metrics for detecting social biases. Ribeiro et al. (2020) perform a test comprising a small set of neutral sentences with simple adjectives and label preserving perturbations to check if the behavior of the LLM differs, and then add a sentiment to the template to check if the model predicts the opposite sentiment Ribeiro et al. (2020). Compared to their work, which focuses on the extent of biased decisions that are made towards protected groups, our work provides insight into decision patterns akin to human cognitive bias where we analyze systematic flaws of language models during a decision-making procedure.
Existing evaluation metrics for societal bias are often based on word embeddings Bolukbasi et al. (2016); Papakyriakopoulos et al. (2020); Viswanath and Zhang (2023), which is not directly applicable for evaluation of decision patterns akin to human cognitive bias. Functional resemblance to cognitive bias is not necessarily embedded in specific tokens but can be reflected in the entire current Tversky and Kahneman (1981) or previous context Echterhoff et al. (2022). To mitigate bias in LLMs, Schick et al. (2021) find that pre-trained language models can recognize corpus-based bias and toxicity. They propose a decoding algorithm that, given a textual description of the behavior, reduces the probability of an LLM producing problematic text.
@@ 2.2 Patterns of Cognitive Bias in LLMs
To address the lack of evaluation metrics for cognitive bias, Lin and Ng (2023) propose metrics for availability and framing bias using dummy inputs and paraphrasing for classification tasks. Our work focuses on generative tasks. Previous work has tackled individual cognitive biases such as representativeness or base rate neglectTalboy and Fuller (2023). However, they evaluate the biases as one-off questions and answers, which require further investigation on generalization. Jones and Steinhardt (2022) test for systematic qualitative errors of LLM responses with human cognitive bias using coding prompts. Cognitive bias can influence many scenarios, but is especially important for high-stakes decisions in human-AI collaboration, as humans might be influenced by the decision assistance that LLMs provide.
@@ 3 Testing for Patterns of Cognitive Bias in LLMs
We classify human-like cognitive bias into three novel categories concerning their manifestation in LLMs. Prompt-based cognitive bias describes a subset of cognitive bias mainly introduced through a user prompt. Inherent bias is instilled in the model through the training data. Inherent bias is not necessarily obvious in the prompt but within the general task itself. Sequential bias is induced by previous model answers.
@@
Anchoring Imagine you are an admissions officer. Typically 30% of students are admitted to the program. Would you admit or reject the following student to the program? The student is from {country} and studied {major} at {school} with a GPA of {gpa}, […]. Your answer was admit. The student is from {country} and studied {major} at {school} with a GPA of {gpa}, […]. What’s your decision?
Status Quo/ Primacy Status Quo:You are aiming to admit a student to your lab. [You have worked with student A during a summer internship.] Here are student features to help your decision: a) [Retain] Student A: They have a background in {background}. They are experienced in {experience}, but need to develop {skill} skills. […] d) Student D: Comes from {university} with strong skills in {skill}. They specialize in {topic}, but lack experience in {experience}. Select option a, b, c or d. Neutral:You are aiming to admit a student to your university lab. Here are student features to help your decision: a) Student A: They have a background in {background}. They are experienced in {experience}, but need to develop {skill} skills. […] Select option a, b, c or d.
@@
We design experiments to test decision patterns similar to human cognitive bias, based on definitions from human cognitive bias literature. Our work aims to align all bias groups (prompt-based, sequential, inherent) as much as possible with the same evaluation metrics. However, the detection of individual biases in each group has to be tackled with separate metrics to be able to account for the nuances of the bias group. For all biases, we strive to find a metric of “consistency”. Compared to evaluating decision patterns on human participants, LLMs have the distinct advantage of being testable under various study conditions through repeated prompting to evaluate consistency. In the following, we describe the creation of the prompt dataset.
@@ 3.1 Sequential Bias
Anchoring bias describes the human tendency to change perception based on an anchor Kahneman et al. (1982). We follow the setup of (Echterhoff et al., 2022), in which decision-makers are influenced (anchored) by their own recent decisions. This setup evaluates bias in sequential setups, compared to one-off prompt-based setups (which we discuss in the next section).
@@
To analyse the influence of previous decisions in language models, we ask the model to take the role of an admissions officer deciding which student to admit to a college study program. We create synthetic student profiles and show them to the language model in a conversation by always adding the previous students and the model’s previous decisions to the context. We perturb different student sets such that the same set of students is exposed to the model in different orders, to observe if LLMs make different decisions for the same students. We show examples of our templates in Table 1.
@@ 3.2 Prompt-Based Cognitive Bias
Status quo bias is a cognitive bias that refers to the tendency of people to prefer and choose the current state of affairs or the existing situation over change or alternative options Samuelson and Zeckhauser (1988). Given a set of questions that differ in their content by providing a default option in the status quo, a biased question can be compared to the same prompt without status quo information (neutral condition). Questions always provide different options to choose from. We take inspiration from Samuelson and Zeckhauser (1988) which biases the user with a status quo option with respect to car brands and investment options to choose from. Given e.g. a current car brand they drive or a current investment, users then have to make a decision to switch their car or investment or keep the status quo.
@@
We develop a template for testing if a model shows decision patterns similar to status quo bias between a neutral question, which has no information on current status, and a status quo question for the student admissions setup. In this case, we ask for a student to be admitted to a research lab given student features, and provide four options to choose from. We define the status quo to be “having worked with student X in a summer internship before”. Our prompting contains no indication of whether working with student X was a good or bad experience beforehand. Other parts of the question and the student options remain the same. From a pool of 16 student profiles, we choose 4 to be displayed at a time and show each student at each position to evaluate if some options are chosen disproportionally.
@@
[…]
@@
To analyze the difference in admissions or rejection behavior, we observe the admissions rate $\frac{1}{n}\sum_{i=0}^{n}d_{i}$ for admission decisions where $d_{i}\in\{0,1\}$ for rejection/admission of a student for all students $i=[0,...,n]$, which should not be affected by the framing of the question.
@@ 3.3 Inherent Cognitive Bias
Bias
# Baseline
Prompts
Factor
Anchoring 5449 ×$\times 1$
Status Quo/Primacy 1008 ×$\times 2$
Framing 1000 ×$\times 3$
Group Attribution 1000 ×$\times 3$
Table 2: Number of baseline prompt instances in our dataset per cognitive bias. For status quo, we provide status quo and non-status quo prompts (hence we have a factor 2). For framing, we provide admit, reject, and neutral framing (factor 3). For group attribution, we provide female, male, and neutral prompts (factor 3). We also provide variations of the prompts for awareness, contrastive, and counterfactual mitigation.
@@ 3.4 BiasBuster Prompt Dataset
In total, we provide a dataset that can be used to test the LLM on patterns akin to human cognitive bias. The dataset consists of 13,465 prompts for the baseline conditions. We show the size of each bias dataset in Table 2. For all our prompts, we use the English language. We publish our dataset on Huggingface.
@@ 4 Mitigating Cognitive Bias in LLMs
There are different approaches to mitigating decision patterns similar to human cognitive bias in LLMs. We group these approaches into zero-shot approaches, which can give additional information about the existence of cognitive bias without giving any examples, few shot approaches which can give examples of specific desired or undesired behavior, and self-mitigation approaches, which use the model to debias themselves (Figure 2).
Figure 2: Overview of different mitigation techniques and comparison to our selfhelp setup, which is tasked to debias its prompts. We give an example of status quo bias, where the bias-inducing part of the prompt (in red) is removed by selfhelp.
@@ 4.2 Few-Shot-Mitigation
For group attribution, we show the same student twice, once as female as male, and ask the model answers to be the same. For framing, we show an example of the same student in different framing and ask the model to give the same admission outcome. For status quo, we show an example where the current student is not the most suitable candidate but is still selected. For anchoring, we show two different orders of the same students with different answers for the individuals (Table 6).
@@
We show examples of counterfactual and contrastive mitigation for each bias in the Appendix in Table 6.